OpenStudy (anonymous):

the sum of two angles is 119 degrees. one angle is 22 degrees lass than twice the other. find the angles.

OpenStudy (anonymous):

x+y = 119 x = 2y-22 x+y = 119 2y - 22 + y = 119 3y = 119+22 y = 141/3 y = 47 x=2y-22 x=2(47)-22 x=94-22 x=72

OpenStudy (anonymous):

x=72 y=47